OpenStudy (anonymous):

what is the arm span of someone who is 5 foot 3 inches 5 foot 1 inches 4 foot 11 inches 5 foot 7 inches 5 foot 6 inches

OpenStudy (anonymous):

arm span has no absolute correlation to height. You could be really tall with small arms or maybe really small with long arms. all people who are 5 foot 7 inches, for example, do not have the same exact length arms

OpenStudy (anonymous):

well as i've heard, the arms span of a person is usually equal to his/ her total height. and yup its a fact !

OpenStudy (anonymous):

Absolutely not. to what degree of accuraccy. "usually equal" does not mean its a fact. it might be approximate. but it certainly isnt absolute

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

well sir you do doubt this fact, you can google it, actually the total arm span of a person is about 3 inches more than that of a person's height, so the answer would be D

OpenStudy (anonymous):

http://www.youtube.com/watch?v=lQGrtVmViw4

OpenStudy (anonymous):

? those arent answer choices. i need the arm span of someone who is 5'3 5'7 ect. if you are not going to answer and help please just go to another question.

OpenStudy (anonymous):

the total arm span of a person= height of a person+3 inches( and please calm down)

OpenStudy (anonymous):

i am calm. but i have askd this question 3 times now. twice on this site and once on another. and i keep getting bad answers like these. i was supposed to turn the project in about 3 hours ago but i couldnt because noone would help me. and since noon is still helping i am still screwed. so please just help.

OpenStudy (anonymous):

what project are you doing? is it on davinci? He is the source of this myth

OpenStudy (anonymous):

In his famous sketch of the "vitruvian man" http://en.wikipedia.org/wiki/Vitruvian_Man

OpenStudy (anonymous):

no, it has nothing to do with davinci

OpenStudy (anonymous):

http://leonardodavinci.stanford.edu/submissions/clabaugh/images/vm/leonardo.jpg you can see there he is trying to show that the armspan and body size make approximate perpendicular radii to the same circle

OpenStudy (anonymous):

well what are you doing then, a psudoscience course? perhaps a module on astrology? maybe even some scientology or intelligent design. sorry i cant help you, good luck though :)

OpenStudy (anonymous):

no, it is on algebra. for example, i am 5 foot 3 inches. my hight in inches is 63 my arm span is 69 inches. i only know my armspan because i meaured myself. but i cannot measure someone i made up

OpenStudy (anonymous):

make them up based on your proportions. so your height is 63 and armspan is 69. \[\frac{69}{63} = 1.095\] right? so generaly \[\frac{armspan}{height} = 1.095\] solve this for armspan (the thing your trying to find) \[armspan = height \times 1.095\] so times the height in inches by 1.095 and you will get the armspan you would have if you were that height. for example i'm 75 inches so my armspan is supposedly\[1.095 \times 75 = 82.142 \] so my armspan is 82.142 (if you were me)

OpenStudy (anonymous):

okay uh cool. so what is the arm span of someone who is 5 foot 1 inches 4 foot 11 inches 5 foot 7 inches 5 foot 6 inches

OpenStudy (anonymous):

aka "do my work for me". No

OpenStudy (anonymous):

-_- screw it, ill just get an F. buh-bye

OpenStudy (anonymous):

http://lmgtfy.com/?q=feet+to+inches then times it by 1.095 its the easiest thing in the world

OpenStudy (anonymous):

you deserve an F if you dont do the work yourself. Thats exactly what F's are for. hhere is another F doing its job.

OpenStudy (anonymous):

this is a very small part considering what else i have to do.

OpenStudy (anonymous):

great, so you wont get an F after all! old mika "-_- screw it, ill just get an F" new mika disagrees "this is a very small part considering what else i have to do." what other questions do you have. are they like this one?